Here is my problem:
Find a number such that if 5 times the number is decreased by 14, the result is twice the opposite of the number.
I get 5X-14= but not sure how to figure out the part of twice the opposite number. Would it be 2x?
Thank you for your help.

Answer by ankor@dixie-net.com(22740)

Find a number such that if 5 times the number is decreased by 14, the result is twice the opposite of the number.
:
The opposite number of x is -x, so you have
5x - 14 = 2(-x)
5x -14 = -2x
5x + 2x = 14
x = 2